### v1

# Center for Advanced Intelligence Project, RIKEN 103-0027, Tokyo, Japan

Abstract
By using the squared slack variables technique, we show that a general polynomial complementarity problem can be formulated as a system of polynomial equations. Thus, the solution set of such a problem is the image of a real algebraic set under a certain projection. This paper points out that, generically, this polynomial system has finitely many complex zeros. In such a case, we use techniques from symbolic computation to compute a univariate representation of the solution set. Consequently, univariate representations of special solutions, such as least-norm and sparse solutions, are obtained. After that, enumerating solutions boils down to solving problems governed by univariate polynomials. We also provide some experiments on small-scale problems with worst-case scenarios. At the end of the paper, we propose a method for computing approximate solutions to copositive polynomial complementarity problems that may have infinitely many solutions.
Keywords:
polynomial complementarity problem solution set least-norm solution sparse solution zero-dimensional ideal univariate representation Gröbner basis Shape Lemma
MSC:
90C33 13P10 13P15 65H04
1 Introduction
Let $f:\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}^{n}$ be a mapping. The complementarity problem, named as in facchinei2003finite , given for $f$ is the problem of finding solutions to the following system:
$x\geq 0,\ f(x)\geq 0,\ \left\langle x\mid f(x)\right\rangle=0,$
where $\langle\alpha\mid\beta\rangle$ is the usual scalar product of two vectors $\alpha,\beta$ in the Euclidean space $\operatorname{\mathbb{R}}^{n}$. If $f$ is an affine mapping, the problem becomes a linear complementarity problem, first introduced by Cottle and Dantzig cottle1968complementary . If $f$ is a real polynomial mapping, i.e. all the components of $f$ are polynomial, then the problem is called a polynomial complementarity problem (PCP, for short) gowda2016polynomial . Moreover, PCP is a special case of a polynomial variational inequality hieu2020solution .
Qualitative properties of PCPs and extension versions have recently been studied; such as existence, stability, and compactness of the solution sets gowda2016polynomial ; hieu2020solution ; zheng2020nonemptiness ; hieu2020notes ; ling2019properties ; wang2020existence ; pham2021complementary , bounds of the solution sets ling2018error ; xu2024bounds ; li2024lower or strict feasibility li2024strict . The method using semidefinite relaxations in zhao2019semidefinite can be modified to enumerate numerically the solutions of a general PCP with finitely many solutions. Computing solutions to a general PCP is a difficult problem, even such a task for problems governed by affine mappings is an NP-hard problem chung1989np .
In the case where $f$ is defined by a tensor, the problem is called a tensor complementarity problem che2016positive ; song2016tensor . Multi-player non-cooperative games, hyper-graph clustering problems, and traffic equilibrium problems can be modeled as tensor complementarity problems huang2019tensor . In the literature, there are several works on solving such tensor problems qi2019tensor by using methods based on systems of tensor equations luo2017sparsest ; xu2019equivalent ; xie2017iterative , potential reduction zhang2019potential , homotopy continuation han2019continuation , smoothing reformulations huang2017formulating , the modulus-based reformulation liu2018tensor , reformulated programming problems luo2017sparsest ; xie2017iterative , reformulated integer programming du2019mixed , and semidefinite relaxations zhao2019semidefinite .
This work is motivated by open problems raised by Huang and Qi in (huang2019tensor, , Problem 1) and (qi2019tensor, , Problem 4) related to using techniques from semi-algebraic geometry to study tensor complementarity problems and the number of solutions in the case where the solution set has finitely many points. Our proposed method addresses these issues for a broad class of problems that includes tensor problems. We are also interested in algorithms computing least-norm and sparse solutions of PCPs; It should be recalled that several algorithms have been proposed for linear or tensor problems, but with special structures chen2016sparse ; xie2017iterative ; luo2017sparsest .
In this paper, we employ the squared slack variables technique, which has been used by many researchers, see e.g. fukuda2017note and references therein, to demonstrate that the solution set of a PCP is the image of a real algebraic set under a certain projection, see Proposition 1. This result indicates that the problem is solvable if and only if the real algebraic set is nonempty.
We introduce a class of PCPs given by polynomial mappings $f$ such that the polynomial systems $\{x_{1}f_{1},\dots,x_{n}f_{n}\}$ have finitely many complex zeros. A PCP belonging to this class is called $\operatorname{\mathcal{D}}_{0}$, and it should be highlighted that this class is almost everywhere, see Proposition 2. Using Gröbner bases, the shape lemma, and results from computational commutative algebra, we prove in Theorem 4.1 the existence of a univariate representation for the solution set of a $\operatorname{\mathcal{D}}_{0}$ problem. Based on the proof, we design an algorithm to compute such a representation, see Algorithm 1. Additionally, algorithms are proposed to compute univariate representations of least-norm or sparse solutions. Once a univariate representation is obtained, we can enumerate the solutions using numerical computation. Furthermore, we discuss the sharp upper bound on the number of solutions of $\operatorname{\mathcal{D}}_{0}$ problems based on the well-known Bézout bounds.
We are also interested in solving PCPs that may not be $\operatorname{\mathcal{D}}_{0}$, whose solution sets are infinite. This paper discusses copositive polynomial mappings and proposes a method to approximate their solutions.
The rest of this paper consists of six sections. Section 2 gives some definitions, notations, and auxiliary results of polynomial complementarity problems and computational commutative algebra. Section 3 gives results concerning the squared slack variables technique. Results on the class of $\operatorname{\mathcal{D}}_{0}$ problems are presented in Section 4 and Section 5. Section 6 is about computing numerical solutions for $\operatorname{\mathcal{D}}_{0}$ problems and gives some experiments. The last section, Section 7, discusses copositive problems.
2 Preliminaries
This section recalls basic notions and results from complementarity problems and computational commutative algebra. Further details can be found in facchinei2003finite and cox2013 .
2.1 Polynomial complementarity problems
Let $\operatorname{\mathbb{R}}$ denote the field of real numbers. The notation $\langle\alpha\mid\beta\rangle$ is used for the usual scalar product of two vectors $\alpha,\beta$ in the Euclidean space $\operatorname{\mathbb{R}}^{n}$, whereas $\|\alpha\|$ denotes the Euclidean norm of $\alpha$. The index set $\{1,\dots,n\}$ is written as $[n]$ for brevity.
Let $f=(f_{1},\dots,f_{n}):\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}^{n}$ be a real polynomial mapping. The polynomial complementarity problem, introduced in gowda2016polynomial , for the function $f$ is the problem of finding a solution $\alpha\in\operatorname{\mathbb{R}}^{n}$ of the following system:
$x\geq 0,\ f(x)\geq 0,\ \left\langle x\mid f(x)\right\rangle=0.$ (2.1)
The problem and its solution set are, respectively, denoted by $\operatorname{CP}(f)$ and $\operatorname{Sol}(f)$. Suppose that $\alpha$ is a solution, so the first two conditions in (2.1) say that $\alpha$ and $f(\alpha)$ belong concurrently to the non-negative orthant of $\operatorname{\mathbb{R}}^{n}$, and the last condition means that $\alpha$ must belong to the surface given by the following equation:
$x_{1}f_{1}(x)+\dots+x_{n}f_{n}(x)=0.$ (2.2)
Remark 1
It is not difficult to see that $\alpha$ is a solution of $\operatorname{CP}(f)$ if and only if $\alpha$ satisfies the following system:
$x\geq 0,\ f(x)\geq 0,\ x_{1}f_{1}(x)=\dots=x_{n}f_{n}(x)=0.$ (2.3)
In this setting, the polynomial equation (2.2) can be written equivalently as a square polynomial system, i.e., as $n$ polynomial equations in $n$ variables. This is the key idea allowing us to apply techniques in solving polynomial equations, especially when the square system has finitely many complex roots.
A least-norm solution of $\operatorname{CP}(f)$ is defined as a solution to the following optimization problem:
$\min\{\|\alpha\|:\alpha\in\operatorname{Sol}(f)\}.$ (2.4)
[…]
One defines a sparse solution of $\operatorname{CP}(f)$ as a solution to the following optimization problem:
$\min\{\|\alpha\|_{0}:\alpha\in\operatorname{Sol}(f)\},$ (2.5)
where $\|\alpha\|_{0}$ is the number of nonzero coordinates of $\alpha$. The set of sparse solutions of $\operatorname{CP}(f)$ is denoted by $\operatorname{Sol_{\operatorname{sp}}}(f)$.
2.2 Zero-dimensional ideals and the Shape Lemma
We denote by $\operatorname{\mathbb{C}}$ the field of complex numbers that is the algebraic closure of the field $\operatorname{\mathbb{R}}$. The ring of all polynomials in variables $x=(x_{1},\dots,x_{n})$ with real coefficients is denoted by $\operatorname{\mathbb{R}}[x]$. An additive subgroup $\operatorname{\mathcal{I}}$ of $\operatorname{\mathbb{R}}[x]$ is said to be an ideal of $\operatorname{\mathbb{R}}[x]$ if $pq\in\operatorname{\mathcal{I}}$ for any $p\in\operatorname{\mathcal{I}}$ and $q\in\operatorname{\mathbb{R}}[x]$. Given $p_{1},\dots,p_{s}$ in $\operatorname{\mathbb{R}}[x]$, we denote by $\left\langle p_{1},\dots,p_{s}\right\rangle$ the ideal generated by $p_{1},\dots,p_{s}$. If $\operatorname{\mathcal{I}}$ is an ideal of $\operatorname{\mathbb{R}}[x]$ then, according to Hilbert’s basis theorem (see, for example Theorem 4 in (cox2013, , Chapter 2, §5)), there exist $p_{1},\dots,p_{s}\in\operatorname{\mathbb{R}}[x]$ such that $\operatorname{\mathcal{I}}=\left\langle p_{1},\dots,p_{s}\right\rangle$.
Let $\operatorname{\mathcal{I}}$ be an ideal of $\operatorname{\mathbb{R}}[x]$. The complex algebraic variety associated to $\operatorname{\mathcal{I}}$ is defined as
$V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}})\coloneqq\{x\in{\operatorname{\mathbb{C}}}^{n}:g(x)=0,\forall g\in\operatorname{\mathcal{I}}\}.$
The real algebraic variety associated to $\operatorname{\mathcal{I}}$ is $V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}})\coloneqq V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}})\cap\operatorname{\mathbb{R}}^{n}$. Recall that the ideal $\operatorname{\mathcal{I}}$ is zero-dimensional if the cardinality of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}})$, denoted by #$\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}})$, is finite. The radical of $\operatorname{\mathcal{I}}$ is the following set:
$\operatorname{rad}(\operatorname{\mathcal{I}})\coloneqq\{p\in\operatorname{\mathbb{R}}[x]:p^{k}\in\operatorname{\mathcal{I}}\text{ for some }k\in\operatorname{\mathbb{N}}\}.$
The set $\operatorname{rad}(\operatorname{\mathcal{I}})$ is also an ideal that contains $\operatorname{\mathcal{I}}$. One says that the ideal $\operatorname{\mathcal{I}}$ is radical if $\operatorname{\mathcal{I}}=\operatorname{rad}(\operatorname{\mathcal{I}})$.
The following lemma asserts that the radical of an ideal is radical; Moreover, the complex (real) varieties of $\operatorname{\mathcal{I}}$ and of its radical coincide, see, for example, Lemma 5 and Theorem 7 in (cox2013, , Chapter 4, §2):
Lemma 1
Let $\operatorname{\mathcal{I}}$ be an ideal of $\operatorname{\mathbb{R}}[x]$. Then, $\operatorname{rad}(\operatorname{\mathcal{I}})$ is radical. Moreover, one has $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}})=V_{\operatorname{\mathbb{C}}}(\operatorname{rad}(\operatorname{\mathcal{I}}))$, hence $V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}})=V_{\operatorname{\mathbb{R}}}(\operatorname{rad}(\operatorname{\mathcal{I}}))$.
Let $<$ be a monomial ordering on $\operatorname{\mathbb{R}}[x]$ and $\operatorname{\mathcal{I}}\neq\{0\}$ be an ideal of $\operatorname{\mathbb{R}}[x]$. We denote by $\operatorname{LT}_{<}(\operatorname{\mathcal{I}})$ the set of all leading terms $\operatorname{LT}_{<}(p)$ for $p\in\operatorname{\mathcal{I}}$, and by $\left\langle\operatorname{LT}_{<}(\operatorname{\mathcal{I}})\right\rangle$ the ideal generated by the elements of $\operatorname{LT}_{<}(\operatorname{\mathcal{I}})$. A subset $G=\{p_{1},\dots,p_{s}\}$ of $\operatorname{\mathcal{I}}$ is said to be a Gröbner basis of $\operatorname{\mathcal{I}}$ with respect to the monomial order $<$ if
$\left\langle\operatorname{LT}_{<}(p_{1}),\dots,\operatorname{LT}_{<}(p_{s})\right\rangle=\left\langle\operatorname{LT}_{<}(\operatorname{\mathcal{I}})\right\rangle.$
Note that a Gröbner basis of $\operatorname{\mathcal{I}}$ really is a generating set of the ideal $\operatorname{\mathcal{I}}$, i.e. $\left\langle p_{1},\dots,p_{s}\right\rangle=\operatorname{\mathcal{I}}$ and every ideal in $\operatorname{\mathbb{R}}[x]$ has a Gröbner basis (see (cox2013, , Cor. 6)). A Gröbner basis $G$ is reduced if the two following conditions hold: i) the leading coefficient of $p$ is $1$, for all $p\in G$; ii) for all $p\in G$, there are no monomials of $p$ lying in $\left\langle\operatorname{LT}_{<}(G\setminus\{p\})\right\rangle$. Every ideal $\operatorname{\mathcal{I}}$ has a unique reduced Gröbner basis. We refer the reader to cox2013 for more details.
Let $\operatorname{\mathcal{I}}$ be a zero-dimensional and radical ideal in $\operatorname{\mathbb{R}}[x]$, and $G$ be its reduced Gröbner basis with respect to the lexicographic ordering with $x_{1}<_{\operatorname{lex}}\cdots<_{\operatorname{lex}}x_{n}$. One says that $\operatorname{\mathcal{I}}$ is in shape position if $G$ has the following shape:
$G=[w,x_{2}-u_{2},\dots,x_{n}-u_{n}],$ (2.6)
where $w,u_{2},\dots,u_{n}$ are polynomials in $\operatorname{\mathbb{R}}[x_{1}]$ and $\deg w=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}})$. When the order $<_{\operatorname{lex}}$ is clear from the context, we omit the subscript $\operatorname{lex}$ in the above notation.
The following lemma, named Shape Lemma, gives us a criterion for the shape position of an ideal.
Lemma 2
(Shape Lemma, gianni89 ) Let $<_{\operatorname{lex}}$ be a lexicographic monomial ordering in $\operatorname{\mathbb{R}}[x]$. Suppose that $\operatorname{\mathcal{I}}$ is a zero-dimensional and radical ideal. If the points in $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}})$ have distinct $x_{1}$-coordinates, then $\operatorname{\mathcal{I}}$ can be made into shape position as in (2.6), where $u_{2},\dots,u_{n}$ are polynomials in $\operatorname{\mathbb{R}}[x_{1}]$ of degrees at most #$\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}})-1$.
3 The squared slack variables technique for general PCPs
Using the squared slack variables technique, we show in this section that the solution set of $\operatorname{CP}(f)$ is the image of a real algebraic set under a certain projection.
Denote by $\operatorname{\mathcal{I}}[f]$ the ideal in $\operatorname{\mathbb{R}}[x]$ generated by the $n$ polynomial products $x_{i}f_{i}$, $i\in[n]$, i.e.
$\operatorname{\mathcal{I}}[f]\coloneqq\left\langle x_{1}f_{1},\dots,x_{n}f_{n}\right\rangle.$ (3.1)
We introduce $2n$ squared slack variables $z=(z_{1},\dots,z_{2n})$ and denote by $\operatorname{\mathcal{I}}[f,z]$ the ideal in $\operatorname{\mathbb{R}}[x,z]$ generated by the $n$ polynomials defining $\operatorname{\mathcal{I}}[f]$, the $n$ quadratic polynomials $z_{i}^{2}-x_{i}$, $i\in[n]$, and the $n$ polynomials $z_{n+i}^{2}-f_{i}$, $i\in[n]$, i.e.,
$\operatorname{\mathcal{I}}[f,z]=\left\langle x_{1}f_{1},\dots,x_{n}f_{n},z_{1}^{2}-x_{1},\dots,z_{n}^{2}-x_{n},z_{n+1}^{2}-f_{1},\dots,z_{2n}^{2}-f_{n}\right\rangle.$ (3.2)
Here and in the following, $\sqrt{c}$ stands for the set of the (two) complex square roots of a complex number $c$, and $\pi_{n}:\operatorname{\mathbb{C}}^{n+m}\to\operatorname{\mathbb{C}}^{n}$, for some non-negative integer $m$, is the projection onto the first $n$ coordinates.
Remark 2
Let $(\alpha,\beta)\in\operatorname{\mathbb{C}}^{n+2n}$ with $\alpha=(\alpha_{1},\dots,\alpha_{n})$. By the definition of $\operatorname{\mathcal{I}}[f,z]$, it is not difficult to see that $(\alpha,\beta)$ is a point in $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$ if and only if $\alpha\in\operatorname{\mathcal{I}}[f]$ and $\beta$ is an element of the following set which is defined by a $2n$-fold Cartesian product:
$\sqrt{\alpha_{1}}\times\dots\times\sqrt{\alpha_{n}}\times\sqrt{f_{1}(\alpha)}\times\dots\times\sqrt{f_{n}(\alpha)}.$ (3.3)
Furthermore, from the fact that #$\#\sqrt{c}$ is finite, for $c\in\operatorname{\mathbb{C}}$, we conclude that the set $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ is finite if any only if $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$ so is.
Proposition 1
One has the following equality:
$\operatorname{Sol}(f)=\pi_{n}\big{(}V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[f,z])\big{)}.$ (3.4)
Proof
Let $\alpha=(\alpha_{1},\dots,\alpha_{n})\in\operatorname{\mathbb{R}}^{n}$ be a point in $\operatorname{Sol}(f)$. Clearly, the coordinates of $\alpha$ satisfy $\alpha_{i}f_{i}(\alpha)=0$, for all $i\in[n]$. Since $\alpha_{i}$ and $f_{i}(\alpha)$ are non-negative, we choose $\hat{z}_{i}\in\sqrt{\alpha_{i}}$ and $\hat{z}_{n+i}\in\sqrt{f_{i}(\alpha)}$ for $i\in[n]$. From (3.2), $(\alpha,\hat{z})$ is a point in the real algebraic set of $\operatorname{\mathcal{I}}[f,z]$. Thus, $\pi_{n}(\alpha,\hat{z})=\alpha$.
To prove the inverse inclusion, we suppose that $\alpha\in\pi_{n}\big{(}V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[f,z])\big{)}$. There exists $\hat{z}\in\operatorname{\mathbb{R}}^{2n}$ such that $(\alpha,\hat{z})\in V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[f,z])$. By the definition of $\operatorname{\mathcal{I}}[f,z]$, one has $\alpha_{i}f_{i}(\alpha)=0$, $\alpha_{i}=\hat{z}_{i}^{2}\geq 0$, and $f_{i}(\alpha)=\hat{z}_{n+i}^{2}\geq 0$, for all $i\in[n]$. Hence, $\alpha$ is a element of $\operatorname{Sol}(f)$.
The equality (3.4) is proved. ∎
Remark 3
From Proposition 1, the problem $\operatorname{CP}(f)$ is solvable if and only if the polynomials defining $\operatorname{\mathcal{I}}[f,z]$ have common real roots. In some computer algebra systems, we can check whether a system of polynomials has real roots. In Maple, for example, we can use $\mathtt{HasRealRoots}$ in the $\mathtt{RootFinding}$ package to perform such tasks for polynomials with rational coefficients.
Example 1
Consider a linear complementarity problem $\operatorname{CP}(f)$ in two variables, where $f(x_{1},x_{2})=(x_{2}-1,x_{1}-1)$. We use four slack variables $(z_{1},\dots,z_{4})$ and consider the following system of polynomials in $\operatorname{\mathbb{R}}[x_{1},x_{2},z_{1},\dots,z_{4}]$:
$\{x_{1}x_{2}-x_{1},\ x_{1}x_{2}-x_{2},z_{1}^{2}-x_{1},z_{2}^{2}-x_{2},z_{3}^{2}-x_{2}+1,z_{4}^{2}-x_{1}+1\}.$
This system has real roots, for example, $(1,1,1,1,0,0)$. We conclude that the complementarity problem has a solution.
Remark 4
The squared slack variables technique has been demonstrated to be an effective method for converting polynomial inequalities into equations. In particular, every basic closed semi-algebraic set $S$ in $\operatorname{\mathbb{R}}^{n}$ given by the inequality constraints $\{g_{1}(x)\geq 0,\dots,g_{m}(x)\geq 0\}$ is the image of the following real algebraic set under projection $\pi_{n}$:
$\{(x,z)\in\operatorname{\mathbb{R}}^{n+m}:z_{1}^{2}-g_{1}(x)=0,\dots,z_{m}^{2}-g_{m}(x)=0\}.$
Consequently, findings on real algebraic sets, e.g., compactness or computing dimension pham2018compactness ; lairez2021computing , can be extended to closed semi-algebraic sets. A limitation of this approach is that the number of variables in the resulting system is $n+m$. However, the degrees of the new variables are only two.
4 Existence of univariate representations for $\operatorname{\mathcal{D}}_{0}$ problems
In this section, we focus on the case where $\operatorname{\mathcal{I}}[f]$ is zero-dimensional, i.e. the complex variety $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ has finitely many points. It should be noted that this condition can be checked by using computer algebra systems.
Definition 1
If the complex variety $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ has finitely many points, we say that the polynomial complementarity problem given by $f$ is a $\operatorname{\mathcal{D}}_{0}$ problem, or $f$ is $\operatorname{\mathcal{D}}_{0}$ for simplicity.
Let $\operatorname{\mathbb{R}}_{d}[x]$ be the set of all polynomials with real coefficients of degree at most $d$. Recall that the dimension of $\operatorname{\mathbb{R}}_{d}[x]$, denoted by $\rho$, is finite and the linear space $\operatorname{\mathbb{R}}_{d}[x]^{n}$ is isomorphic to $\operatorname{\mathbb{R}}^{n\times\rho}$. For every $f=(f_{1},\dots,f_{n})$ in $\operatorname{\mathbb{R}}_{d}[x]^{n}$, $f_{i}$ can be written (uniquely) with coefficients $a_{i}\in\operatorname{\mathbb{R}},Q_{i}\in\operatorname{\mathbb{R}}^{\rho-1}$ as follows: $f_{i}(x)=a_{i}+Q_{i}X^{T}$, where
$X=(x_{1},\dots,x_{n},x_{1}^{2},x_{1}x_{2},\dots,x_{n}^{2},\dots,x_{n}^{d})^{T}.$ (4.1)
Recall that a property is said to be generic on $\operatorname{\mathbb{R}}_{d}[x]^{n}$ if it holds on a subset $\mathcal{O}\subset\operatorname{\mathbb{R}}_{d}[x]^{n}$ such that its complement $\operatorname{\mathbb{R}}_{d}[x]^{n}\setminus\mathcal{O}$ has Lebesgue measure zero.
The following proposition relates to the genericity of $\operatorname{\mathcal{D}}_{0}$ whose proof is postponed to the Appendix.
Proposition 2
The property of being $\operatorname{\mathcal{D}}_{0}$ is generic in $\operatorname{\mathbb{R}}_{d}[x]^{n}$.
4.1 A sharp bound on the cardinality of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$
It should be highlighted that if we use $m=2n$ squared slack variables as in Remark 3, the cardinality of the complex algebraic set can increase $2^{2n}$ times. However, in the setting of the complementarity problem, the multiplier is reduced from $2^{2n}$ to $2^{n}$, as in the following proposition. Recall that the degree of the polynomial mapping $f$ is the following number:
$d:=\max\{\deg f_{i}:i\in[n]\}.$
Proposition 3
Let $f$ be $\operatorname{\mathcal{D}}_{0}$. Then, the following inequalities hold:
[…]
Proof
We prove that for a given solution $\alpha$ the cardinality of the set (3.3) does not exceed $2^{n}$. Indeed, for each pair $\{\alpha_{i},f_{i}(\alpha)\}$, there is at least one entry that is zero, so the cardinality of $\sqrt{\alpha_{i}}\times\sqrt{f_{i}(\alpha)}$ is at most two. Hence, the cardinality of the set (3.3) does not exceed $2^{n}$. This implies that
[…]
Applying Bézout’s theorem (cox2013, , Chapter 8, §7) to the system defining $\operatorname{\mathcal{I}}[f]$, one has #$\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\leq(d+1)^{n}$. Therefore, the second inequality in (4.2) is obtained. ∎
We are also interested in the maximum number of solutions of $\operatorname{\mathcal{D}}_{0}$ problems with fixed degree $d$ and number of variables $n$. A sharp upper bound on this number is given in the following remark.
Remark 5
Let $d$ be fixed. Then the maximum number of solutions of $\operatorname{CP}(f)$, where $f$ is of degree $d$ and is $\operatorname{\mathcal{D}}_{0}$, does not exceed $(d+1)^{n}$. Indeed, it is clear that #$\#\operatorname{Sol}(f)\leq\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\leq(d+1)^{n}$, where the last number is the Bézout bound of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$.
Moreover, if $d\in\operatorname{\mathbb{N}}$ is even, then the maximum is reached, i.e., there exists $p\in\operatorname{\mathbb{R}}[x]^{n}$ of degree $d$ and being $\operatorname{\mathcal{D}}_{0}$ such that #$\#\operatorname{Sol}(p)=(d+1)^{n}$. We consider the polynomial map $p$ given by $p_{i}(x)=\prod_{j=1}^{d}(x_{i}-j),\ i\in[n]$. Clearly, $x_{i}p_{i}$ has $d+1$ non-negative (real) roots, listed by $\{0,1,\dots,d\}$. It is not difficult to check that
$V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[p])=V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[p])=\big{\{}(a_{1},\dots,a_{n})\in\operatorname{\mathbb{R}}^{n}:a_{i}\in\{0,1,\dots,d\}\big{\}}$,
and that the solution set $\operatorname{Sol}(p)$ coincides with $V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[p])$. Hence, the number of points in $\operatorname{Sol}(p)$ is $(d+1)^{n}$.
Remark 6
It should be noted that the number of variables is generally the most important factor in the complexity of symbolic computation on polynomials. We remark here about a technique for reducing the number of slack variables. It is logical to replace $x_{i}$, which is required to be non-negative, by $y_{i}^{2}$ in $f(x)$, $i\in[n]$. Then, the system (2.3) becomes the following system:
$f(y^{2})\geq 0,\ y_{1}f_{1}(y^{2})=0,\dots,y_{n}f_{n}(y^{2})=0,$
where $y^{2}:=(y_{1}^{2},\dots,y_{n}^{2})$. We can add $n$ squared slack variables to the above system (regarding the condition $f(y^{2})\geq 0$), then we get a new system of $2n$ variables. However, the degree of this system could be $12d+$ 1, then the cardinality of the complex variety can reach $2^{n}(2d+1)^{n}$. This number is larger than the bound $2^{n}(d+1)^{n}$ given in (4.2).
4.2 Univariate representations of the solutions
Theorem 4.1
Let $f$ be $\operatorname{\mathcal{D}}_{0}$. Then, there exist polynomials $w,v_{1},\dots,v_{n}$ in $\operatorname{\mathbb{R}}[t]$, with $\deg v_{i}<\deg w$, $i\in[n]$, and $\deg w=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$, such that
$\operatorname{Sol}(f)=\big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:w(t)=0,t\in\operatorname{\mathbb{R}}\big{\}}.$ (4.3)
Proof
Since $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ is finite, according to (rouillier1999, , Lemma 2.1), there exists an invertible $3n\times 3n$-matrix $H$ with real entries such that, if we change the variables $(x,z)=Hy$ into the following polynomials
$x_{1}f_{1},\dots,x_{n}f_{n},z_{1}^{2}-x_{1},\dots,z_{n}^{2}-x_{n},z_{n+1}^{2}-f_{1},\dots,z_{2n}^{2}-f_{n},$ (4.4)
obtain respectively polynomials $h_{1}(y),\dots,h_{3n}(y)$, and denote by the ideal in $\operatorname{\mathbb{R}}[y_{1},\dots,y_{3n}]$
$\operatorname{\mathcal{I}}_{h}\coloneqq\left\langle h_{1},\dots,h_{3n}\right\rangle,$
then the points in $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}_{h})$ have distinct $y_{1}$-coordinates. Thanks to Lemma 1, we have $V_{\operatorname{\mathbb{C}}}(\operatorname{rad}(\operatorname{\mathcal{I}}_{h}))=V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}_{h})$. In view of the assumption and Proposition 3, $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is zero-dimensional and radical.
Clearly, the ideal $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ satisfies the condition of the Shape Lemma. So, there exist $3n$ polynomials in $\operatorname{\mathbb{R}}[y_{1}]$, says $w$, $u_{2},\dots,u_{3n}$, with $\deg w=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$ and $\deg u_{i}<\deg w$, $i=2,\dots,3n$, and such that $[w,y_{2}-u_{2},\dots,y_{3n}-u_{3n}]$ is the reduced Gröbner basis with respect to the lexicographic order with $y_{1}<\cdots<y_{3n}$ of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$.
We write $u(t):=(t,u_{2}(t),\dots,u_{3n}(t))$ for short. From the change of variables, one has
$V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])=\big{\{}H(u(t))^{T}:w(t)=0\big{\}}.$
Recall that the entries of $H$ are real, then that ones of $H^{-1}$ so are. Suppose that $\tau$ is a root of $w(t)$. If $\tau$ is real, then all coordinates of $H(u(\tau))^{T}$ are real, hence this vector belongs to $V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[f,z])$. Conversely, if $H(u(\tau))^{T}$ in $\operatorname{\mathbb{R}}^{3n}$ then
$u(\tau)^{T}=\left(H^{-1}H\right)(u(\tau))^{T}\in\operatorname{\mathbb{R}}^{3n}.$
Hence, $\tau$ belongs to $\operatorname{\mathbb{R}}$. We conclude that the real algebraic variety $\operatorname{\mathcal{I}}[f,z]$ is given by
$V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[f,z])=\big{\{}H(u(t))^{T}:w(t)=0,t\in\operatorname{\mathbb{R}}\big{\}}.$ (4.5)
Let $v_{1}(t),\dots,v_{n}(t)$ be the first $n$ coordinates of $H(u(t))$. Since polynomials $u_{2},\dots,u_{3n}$ have real coefficients, so do $v_{1},\dots,v_{n}$. Moreover, if $\deg v_{i}\geq\deg w$ then we replace $v_{i}$ by the remainder of the division of $v_{i}$ by $w$, for $i\in[n]$.
The conclusion of the theorem follows from (4.5) and (3.4). ∎
The system $[w,v_{1},\dots,v_{n}]$ mentioned in Theorem 4.1 is univariate and such a system is called to as a univariate representation of the solutions of $\operatorname{CP}(f)$.
To illustrate Theorem 4.1 and its proof, we consider the following example.
Example 2
We consider the complementarity problem given in Example 1. From (3.1), the ideal $\operatorname{\mathcal{I}}[f]$ is given by $\operatorname{\mathcal{I}}[f]=\left\langle x_{1}x_{2}-x_{1},\ x_{1}x_{2}-x_{2}\right\rangle$. This ideal is zero-dimensional as its complex variety contains two distinct points, $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])=\left\{(0,0),(1,1)\right\}$. We use four slack variables $z=(z_{1},\dots,z_{4})$ and consider the following ideal in $\operatorname{\mathbb{R}}[x,z]$:
$\operatorname{\mathcal{I}}[f,z]=\left\langle x_{1}x_{2}-x_{1},\ x_{1}x_{2}-x_{2},z_{1}^{2}-x_{1},z_{2}^{2}-x_{2},z_{3}^{2}-x_{2}+1,z_{4}^{2}-x_{1}+1\right\rangle.$
By changing variables $(x,z)=Hy$, where $H$ is a $6\times 6$-matrix as follows:
$H:=\begin{bmatrix}1&2&\cdots&6\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\end{bmatrix},$
we obtain six polynomials $h_{1},\dots,h_{6}$ in variables $y$. The ideal $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position with the graded lexicographic order with $y_{1}<y_{2}<\cdots<y_{6}$, with the Gröbner basis of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$, say $[w(y_{1}),y_{2}-u_{2}(y_{1}),\dots,y_{6}-u_{6}(y_{1})]$. The first two coordinates of $H(t,u_{2},\dots,u_{6})$ are denoted by $v_{1},v_{2}$. So we obtain $w$ and $v_{1},v_{2}$ as follows:
$w(t)=t^{8}+4t^{7}+78t^{6}+392t^{5}-5247t^{4}-11228t^{3}-5324t^{2}-11616t$, $v_{1}(t)=(-395708t^{7}-1758766t^{6}-31180792t^{5}-167371327t^{4}+2037780580t^{3}+5545360564t^{2}+2068565664t+5710973125)/5710973125$, $v_{2}(t)=(395708t^{7}+1758766t^{6}+31180792t^{5}+167371327t^{4}-2037780580t^{3}-5545360564t^{2}-2068565664t-5710973125)/5710973125$.
Polynomial $w$ has eight roots including $\{-8,-2,0,6,i,-i,11i,-11i\}$. According to (4.3), the solution set is given by
$\operatorname{Sol}(f)=\big{\{}(v_{1}(t),v_{2}(t)),t\in\{-8,-2,0,6\}\big{\}}=\{(1,1)\}.$
4.3 Univariate representations of least-norm solutions and sparse solutions
Once a univariate representation for the solutions of a $\operatorname{\mathcal{D}}_{0}$ problem is obtained, we can provide a univariate representation for the set of least-norm solutions $\operatorname{Sol_{\operatorname{ln}}}(f)$. In particular, we have the following theorem:
Theorem 4.2
Let $f$ be $\operatorname{\mathcal{D}}_{0}$ and $[w,v_{1},\dots,v_{n}]$ be a univariate representation of $\operatorname{Sol}(f)$. Then, there exists $\varphi$ in $\operatorname{\mathbb{R}}[t]$ such that
$\operatorname{Sol_{\operatorname{ln}}}(f)=\big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:t\in\arg\min\{\varphi(t):w(t)=0,t\in\operatorname{\mathbb{R}}\}\big{\}}.$ (4.6)
Proof
It is clear that the following real function
$\hat{\varphi}(t):=v_{1}^{2}(t)+\dots+v_{n}^{2}(t),\ t\in\operatorname{\mathbb{R}},$ (4.7)
is a univariate polynomial with real coefficients. Denote by $\varphi$ the remainder of the division of $\hat{\varphi}$ by $w$. One has $\deg\varphi<\deg w$. Let $\gamma$ be the minimum of $\varphi(t)$ over the finite set $\{t\in\operatorname{\mathbb{R}},w(t)=0\}$. It is not difficult to prove that $\hat{\alpha}$ is a least-norm solution of $\operatorname{CP}(f)$ if and only if there exists a real root $\tau$ of $w$ such that $\varphi(\tau)=\gamma$ and $\hat{\alpha}=(v_{1}^{2}(\tau),\dots,v_{n}^{2}(\tau))$. Hence, (4.6) is proved. ∎
[…]
Let $[w,v_{1},\dots,v_{n}]$ be a univariate representation of $\operatorname{Sol}(f)$ and $\alpha$ be a solution of $\operatorname{CP}(f)$. Let $\ell$ be a subset of $[n]$. We denote by ¯$\bar{\ell}=[n]\setminus\ell$ and $w_{\ell}$ by the greatest common divisor defined as follows:
$w_{\ell}:=\gcd(w,v_{i},i\in\ell).$
Theorem 4.3
Let $f$ be $\operatorname{\mathcal{D}}_{0}$ and $[w,v_{1},\dots,v_{n}]$ be a univariate representation of $\operatorname{Sol}(f)$. Suppose that sparse solutions have $k$ zero coordinates. Then
$\operatorname{Sol_{\operatorname{sp}}}(f)=\bigcup_{\ell\subseteq[n],k=\#\ell}\Big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:w_{\ell}(t)=0,w_{\bar{\ell}}(t)\neq 0,t\in\operatorname{\mathbb{R}}\Big{\}}.$ (4.8)
Proof
Let $\alpha$ be a sparse solution. Suppose that $\ell$ is the index set of the zero coordinates of $\alpha$. So, there are some real roots $\tau$ of $w$ such that $\alpha=\big{(}v_{1}(\tau),\dots,v_{n}(\tau)\big{)}$ and $v_{i}(\tau)=0$ for $i\in\ell$, and the rest of the coordinates $v_{i}(\tau)\neq 0$ for $i\in\bar{\ell}$. Hence $\tau$ is a real root of $w_{\ell}$ and $\tau$ is not a real root of $w_{\bar{\ell}}$. Therefore, the inclusion $\subseteq$ of (4.8) is proved. The converse inclusion is similarly demonstrated. ∎
5 Algorithms computing univariate representations
Based on the proof of Theorem 4.1, we first design an algorithm (Algorithm 1) to compute a univariate representation of the solution set of a $\operatorname{\mathcal{D}}_{0}$ problem. We then briefly describe algorithms for computing univariate representations for least-norm and sparse solutions based on Algorithm 1.
5.1 Computing univariate representations of $\operatorname{Sol}(f)$
Input and output. The input of Algorithm 1 is a polynomial mapping $f$ being $\operatorname{\mathcal{D}}_{0}$. The output includes a list of univariate polynomials $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ such that $\operatorname{Sol}(f)$ can be represented through the univariate polynomials as in (5.1).
Algorithm 1 Computing a univariate representation of $\operatorname{Sol}(f)$
Input: $f\in\operatorname{\mathbb{R}}[x]^{n}$ such that $f$ is $\operatorname{\mathcal{D}}_{0}$
Output: $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$
$\operatorname{Sol}(f)=\Big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:w(t)=0,t\in\operatorname{\mathbb{R}}\Big{\}},$ (5.1)
where $\deg w=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$ and $\deg v_{i}<\deg w$, $i\in[n]$
  • 1:
    Set $s\coloneqq 0$ and $\operatorname{\mathtt{shape}}\coloneqq$ false
  • 2:
    While $\operatorname{\mathtt{shape}}=$ false do
    • 3:
      Change variables $(x,z)=Hy$ in the polynomials in (4.4), respectively, obtain polynomials denoted by $g_{i}(y)$, $i\in[3n]$, where
      $H:=\begin{bmatrix}1&s&s^{2}&\cdots&s^{3n-1}\\ 0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\end{bmatrix}^{-1}\in\operatorname{\mathbb{R}}^{3n\times 3n}$ (5.2)
    • 4:
      Set $\operatorname{\mathcal{I}}_{h}\coloneqq\left\langle h_{1},\dots,h_{3n}\right\rangle$ in $\operatorname{\mathbb{R}}[y_{1},\dots,y_{3n}]$ and compute the reduced Gröbner basis $G$ of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ with respect to the lexicographic order $y_{1}<y_{2}<\cdots<y_{3n}$
    • 5:
      If the basis $G$ has the following form
      $G=[\hat{w},y_{2}-u_{2},\dots,y_{3n}-u_{3n}],$ (5.3)
      where $\hat{w},u_{2},\dots,u_{3n}$ are in $\operatorname{\mathbb{R}}[y_{1}]$, i.e. $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position, then set $\operatorname{\mathtt{shape}}\coloneqq$ true, else set $s\coloneqq s+1$
  • 6:
    Define by $(v_{1},\dots,v_{n})$ the first $n$ coordinates of $H(t,u_{2},\dots,u_{3n})$
  • 7:
    Return $[w,v_{1},\dots,v_{n}]$
Description. At the beginning, we set $s=0$ and $\operatorname{\mathtt{shape}}=$ false. The variable $\operatorname{\mathtt{shape}}$ is a boolean. The while loop at Line 2 will stop if $\operatorname{\mathtt{shape}}$ is true. Within the while loop, at Line 3, with $s$ given, we define matrix $H$ depending on $s$ as in (5.2) and change the variables $(x,z)=Hy$ in the polynomials in (4.4) and obtain $h_{i}$, $i\in[3n]$. At Line 4, we set
$\operatorname{\mathcal{I}}_{h}\coloneqq\left\langle h_{1},\dots,h_{3n}\right\rangle\subset\operatorname{\mathbb{R}}[y_{1},\dots,y_{3n}]$
and compute the reduced Gröbner basis $G$ of the radical of the ideal $\operatorname{\mathcal{I}}_{h}$, that is $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$, with respect to the lexicographic order with $y_{1}<y_{2}<\cdots<y_{3n}$. At Line 5, we set the value of $\operatorname{\mathtt{shape}}$ as true if $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position otherwise, we increase the value $s\coloneqq s+1$.
When the while loop ends, the reduced Gröbner basis $G$ has the form (5.3), then $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position with
$\operatorname{rad}\left(\operatorname{\mathcal{I}}_{h}\right)=\left\langle w,y_{2}-u_{2},\dots,y_{3n}-u_{3n}\right\rangle.$
The correctness of Algorithm 1 is proved in the following proposition.
Theorem 5.1
Let $f$ be $\operatorname{\mathcal{D}}_{0}$. On input $f$, Algorithm 1 terminates and returns a univariate representation of the solution set $\operatorname{Sol}(f)$.
Proof
Denote by $\delta$ the cardinality of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$, according to (rouillier1999, , Lemma 2.1), there is a number $j$ in the set $\{0,\dots,(3n-1)\delta(\delta-1)/2\}$ such that the linear function $u\coloneqq\left\langle(x,z)\mid(1,s,\dots,s^{3n-1})\right\rangle$ separates $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$, i.e., $u(a)\neq u(b)$ for any distinct points $a,b$ in $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$. After changing the variables at Line 3, namely, $(x,z)=Hy$, the radical of $\operatorname{\mathcal{I}}_{h}$ satisfies the assumptions of Lemma 2. Ideal $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ is in shape position and the while loop ends within at most $(3n-1)\delta(\delta-1)/2+1$ iterations.
After the while loop, the reduced Gröbner basis of $\operatorname{rad}(\operatorname{\mathcal{I}}_{h})$ written as at Line 5, $G=[w,y_{2}-u_{2},\dots,y_{n}-u_{3n}]$. After Line 6, one has
$\big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:w(t)=0,t\in\operatorname{\mathbb{R}}\big{\}}=\pi_{n}(V_{\operatorname{\mathbb{R}}}(\operatorname{\mathcal{I}}[f,z])).$
From the latter equality and (1), we get (5.1). Therefore, we claim that the output of Algorithm 1 provides a univariate representation of $\operatorname{CP}(f)$. ∎
Remark 7
As seen in the proof of Theorem 5.1, the while loop is executed at most $(3n-1)\delta(\delta-1)/2+1$ times, where $\delta=\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$. By Proposition 3 we get an upper bound on the number of iterations of the loop by
$(3n-1)2^{n}(d+1)^{n}(2^{n}(d+1)^{n}-1)/2+1.$
However, in practice the matrix $H$ should be chosen randomly with rational entries, hence we do not need to use the while loop at Line 2.
5.2 Computing univariate representations of $\operatorname{Sol_{\operatorname{ln}}}(f)$
A univariate representation of $\operatorname{Sol_{\operatorname{ln}}}(f)$ can be computed based on the previous algorithm. To complete this task, we propose Algorithm 2, but here we briefly describe it and skip the proof of its correctness. Once a univariate representation $[w,v_{1},\dots,v_{n}]$ of $\operatorname{Sol}(f)$ obtained as the output of Algorithm 1, as mentioned in the proof of Theorem 4.2, we compute $\hat{\varphi}$ as in (4.7) and then $\varphi$ is the remainder of the division of $\hat{\varphi}$ by $w$.
Algorithm 2 Computing a univariate representation of $\operatorname{Sol_{\operatorname{ln}}}(f)$
Input: $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$, where $f$ is $\operatorname{\mathcal{D}}_{0}$
Output: $[w,v_{1},\dots,v_{n},\varphi]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol_{\operatorname{ln}}}(f)$ satisfying (4.6), i.e.,
$\operatorname{Sol_{\operatorname{ln}}}(f)=\big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:t\in\arg\min\{\varphi(t):w(t)=0,t\in\operatorname{\mathbb{R}}\}\big{\}}$
  • 1:
    Compute $\hat{\varphi}(t)=v_{1}^{2}(t)+\dots+v_{n}^{2}(t)$ and $\varphi$ the remainder of the division of $\hat{\varphi}$ by $w$
  • 2:
    Return $[w,v_{1},\dots,v_{n},\varphi]$
5.3 Computing univariate representations of $\operatorname{Sol_{\operatorname{sp}}}(f)$
We design Algorithm 3 to compute a univariate representation of the sparse solution set of a $\operatorname{\mathcal{D}}_{0}$ problem. The output of Algorithm 3 consists of a subset $\Omega\subset 2^{[n]}$ and $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ such that, for every element $\ell$ in $\Omega$, the following set is nonempty:
$\{t\in\operatorname{\mathbb{R}}:w_{\ell}(t)=0,w_{\bar{\ell}}(t)\neq 0\},$ (5.4)
and for each real number $\tau$ in the latter set, $(v_{1}(\tau),\dots,v_{n}(t))$ is a sparse solution whose coordinates at $\ell$ are zero.
Algorithm 3 Computing a univariate representation of $\operatorname{Sol_{\operatorname{sp}}}(f)$
Input: $[w,v_{1},\dots,v_{n}]\subset\operatorname{\mathbb{R}}[t]$ a univariate representation of $\operatorname{Sol}(f)$, where $f$ is $\operatorname{\mathcal{D}}_{0}$
Output: $\Omega$ and $[w,v_{1},\dots,v_{n}]$ such that
$\operatorname{Sol_{\operatorname{sp}}}(f)=\bigcup_{\ell\in\Omega}\Big{\{}\big{(}v_{1}(t),\dots,v_{n}(t)\big{)}:\{t\in\operatorname{\mathbb{R}}:w_{\ell}(t)=0,w_{\bar{\ell}}(t)\neq 0\}\neq\emptyset\Big{\}}.$ (5.5)
  • 1:
    Set $k:=0$ and $\Omega:=\emptyset$
  • 2:
    For $j$ from $n$ to $1$ do
    • 3:
      For each $\ell\subseteq[n]$ such that #$\#\ell=j$, if $\{t\in\operatorname{\mathbb{R}}:w_{\ell}(t)=0,w_{\bar{\ell}}(t)\neq 0\}\neq\emptyset$, then set $k:=j$ and $\Omega:=\Omega\cup\ell$
    • 4:
      If $k>0$, then stop the for loop at Line 2
  • 5:
    Return $\Omega$ and $[w,v_{1},\dots,v_{n}]$
Description. In Algorithm 3, $k$ indicates the number of zero coordinates of a sparse solution, and $\Omega\subset 2^{[n]}$ is displayed as the set of all $\ell\subset[n]$ with #$\#\ell=k$ such that there is a sparse solution whose coordinates at $\ell$ are zero. In the initial stage, we set $k=0$ and $\Omega=\emptyset$, which indicates that the problem lacks sparse solutions.
Since $k$ belongs to $\{0,1,\dots,n\}$ and is the maximum number of zero coordinates, in Line 2 we reduce the value of $j$ starting from $n$. The for loop is stopped when the first index set $\ell$, of a sparse solution, is found. We note that, for each $j$, there are $n!/(j!(n-j)!)$ subset of $[n]$ with cardinality $j$. Once a sparse solution is found, in line 3, for each $\ell\subseteq[n]$ with $k$ elements, if $w_{\ell}$ has real roots and $w_{\bar{\ell}}$ has no real roots, then we let the index set to belong to $\Omega$.
Correctness. The correctness of Algorithm 3 is briefly explained as follows: After the for loop in Line 2, $k$ is the value of $\|\alpha\|_{0}$ for a sparse solution $\alpha$. Clearly, $\ell\in\Omega$ if and only if the set (5.4) is non-empty. Furthermore for each $\tau$ in the set (5.4), $(v_{1}(\tau),\dots,v_{n}(\tau))$ is a sparse solution whose coordinates at $\ell$ are zero. Hence, we get the equality (5.5).
Remark 8
The Shape Lemma still holds if we replace the field of real numbers $\operatorname{\mathbb{R}}$ with the field of rational numbers $\operatorname{\mathbb{Q}}$, i.e. if $\operatorname{\mathcal{I}}$ is an ideal in the ring $\operatorname{\mathbb{Q}}[x]$ and satisfies the conditions in Lemma 2 then the involved polynomials $w,u_{2},\dots,u_{n}$ have rational coefficients. Therefore, if the input polynomial mapping $f$ has rational coefficients, then the univariate polynomials in the aforementioned univariate representations also have rational coefficients.
6 Computing numerical solutions for $\operatorname{\mathcal{D}}_{0}$ problems
This section deals with enumerating (least-norm, sparse) solutions, i.e. explicitly computing all such solutions, of a polynomial complementarity problem $\operatorname{CP}(f)$, where $f$ is $\operatorname{\mathcal{D}}_{0}$.
6.1 Enumerating solutions
It is worth to remark that the number of real roots of $w$ can be greater than the cardinality of $\operatorname{Sol}(f)$; in particular, from the proof of Proposition 3, one has
#$\#\{t\in\operatorname{\mathbb{R}}:w(t)=0\}\leq 2^{n}\times\#\operatorname{Sol}(f).$
The equality can be reached. For example, in Example 1, $w$ has four real roots, says $-8,-2,0,6$, while the problem has only one solution $(1,1)$. Due to the one-to-many correspondence, enumerating $\operatorname{Sol}(f)$ by directly approximating the real roots of $w$ may be incorrect. The real roots can be computed numerically; by evaluating the values of $v_{1},v_{2}$ at these points, we obtain four approximate solutions of $\operatorname{CP}(f)$ as follows:
(1.0000012030, 0.9999997855), (1.0000000063, 1.0000000010),
(1.0000000000, 1.0000000000), (0.9999998072, 1.0000000130),
which are different from each other, but very close to the exact solution $(1,1)$.
From the above issue, in order to enumerate the points in $\operatorname{Sol}(f)$ numerically, we propose the following procedure using a given precision pair $\gamma=(\gamma_{1},\gamma_{2})$, where $\gamma_{1}>0$ and $\gamma_{2}>0$, including two phases as follows:
  • Phase 1. We first compute a univariate representation $[w,v_{1},\dots,v_{n}]$ of $\operatorname{Sol}(f)$ by using Algorithm 1. With the precision $\gamma_{1}$, we next compute numerically the real roots of $w$ and denote by
    $A=\big{\{}\big{(}v_{1}(\tau),\dots,v_{n}(\tau)\big{)}:\tau\text{ is a real roots of }w\text{ with precision }\gamma_{1}\big{\}}.$
For each exact solution $\alpha$, there can be $2^{n}$ points in $A$, denoted by the set $A_{\alpha}$, which are close to $\alpha$.
  • Phase 2. We then reduce $A_{\alpha}$ to one point with precision $\gamma_{2}$ using the following criterion: if $\alpha^{1}$ and $\alpha^{2}$ are in $A$ and $\|\alpha^{1}-\alpha^{2}\|<\gamma_{2}$, then we remove one point of them from the set $A$.
Going back to Example 1, we take $\gamma_{1}=10^{-6}$ and $\gamma_{2}=10^{-4}$, after Phase 2, the four approximate solutions are reduced to $(1,1)$.
Remark 9
The method using semidefinite relaxations in zhao2019semidefinite can be modified to enumerate the solutions of $\operatorname{\mathcal{D}}_{0}$ problems, since they have finitely many solutions. Our method in this paper does this task without using semidefinite relaxations; this method uses the Shape Lemma employed in recent polynomial optimization work, e.g., magron2023 ; hieu2023computing . A method using Gröbner bases to compute solutions to such problems has been briefly described in yang1999grobner , but the system presented there still contains inequalities. Furthermore, the state-of-the-art algorithms for computing real roots of polynomial systems having finitely many complex zeros by computer algebra methods are discussed in berthomieu .
6.2 Experiments on small-scale problems with worst-case scenarios
It should be noted that complementarity problems governed by nonlinear or tensor/polynomial data (without a special structure) are generally hard to solve because of the algorithmic difficulties that arise with higher degrees. Thus, the sizes solved in other papers, e.g. (zhao2019semidefinite, , Section 4), are small. The present paper also faces the same issue.
To see the computational behavior of the procedure proposed in Section 6.1, we report our experiments on small-scale problems with rational coefficients such that their numbers of solutions are maximal.
The procedure is implemented in Maple, and the results are obtained on an Intel i7 - 8665U CPU (1.90GHz) with 32 GB of RAM. The package $\mathtt{Groebner}$ can handle symbolic tasks in Algorithm 1. To compute the radical of an ideal, we use the command $\mathtt{Radical}$; to compute the reduced Gröbner basis of an ideal with respect to a given lexicographic order, we use the command $\mathtt{Basis}$. To enumerate solutions, we use numerical computations by isolating real roots of $w$. We set the variable $\texttt{Digits}:=10+\deg w$, where Digits controls the number of digits that Maple uses when making calculations with software floating-point numbers, and the second accuracy $\gamma_{2}$ is fixed $10^{-4}$.
We perform our experiment with two families of small-scale complementarity problems. The first family is given by the polynomials in Remark 5 with $n=2$, i.e. $p=(p_{1},p_{2})$, where $p_{i}=(x_{i}-1)(x_{i}-2)\dots(x_{i}-d)$ for $i=1,2$. Let $d$ be in $\{2,4,6\}$, the computational results of the procedure described in Section 6.1 are given in the following table:
Alg 1 all solutions ln solutions sp solutions
$d$ $\deg w$ time #$\#\operatorname{Sol}$ time #$\#\operatorname{Sol_{\operatorname{ln}}}$ time #$\#\operatorname{Sol_{\operatorname{sp}}}$ time
2 36 0.59 9 9.40 1 0.80 1 1.10
4 100 6.95 25 80.8 1 11.7 1 4.35
6 144 119 49 616 1 98.3 1 53.7
Table 1. Computational results on $p$
In Table 1, the time is given in seconds; #$\#\operatorname{Sol}$, #$\#\operatorname{Sol_{\operatorname{ln}}}$, and #$\#\operatorname{Sol_{\operatorname{sp}}}$ are respectively the cardinality of the solution set, least-norm solution set, and sparse solution set of $\operatorname{CP}(p)$.
The second family is quadratic complementarity problems, i.e. $d=2$ is fixed, given $q=(q_{1},\dots,q_{n})$, where $q_{i}=x_{i}^{2}-3x_{i}+1$ for $i\in[n]$. Let $n$ run from $1$ to $3$, the computational results are given in the following table:
Alg 1 all solutions ln solutions sp solutions
$n$ $\deg w$ time #$\#\operatorname{Sol}$ time #$\#\operatorname{Sol_{\operatorname{ln}}}$ time #$\#\operatorname{Sol_{\operatorname{sp}}}$ time
1 6 0.12 3 0.44 1 0.14 1 0.26
2 36 0.65 9 9.65 1 0.79 1 1.13
3 216 126 27 1084 1 120 1 4474
Table 2. Computational results on $q$
In Table 2, for the case $n=3$, the time to compute the sparse solution set is very expensive because of the computation of $w_{[3]}$ the greatest common divisor of $w,v_{1},v_{2}$, and $v_{3}$ whose degrees are 216 or 215 with large-scale coefficients; the time for only this task is $4431$ seconds.
7 Solving copositive PCPs
This section is devoted to solving another class of PCPs that may not be $\operatorname{\mathcal{D}}_{0}$, i.e., the solution set may be infinite. In particular, we focus on PCPs where involved mappings are copositive and approximate them by $\operatorname{\mathcal{D}}_{0}$ problems.
One says that a polynomial mapping $g:\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}^{n}$ is copositive if $\left\langle\alpha\mid g(\alpha)\right\rangle\geq 0$ for all $\alpha$ in $\operatorname{\mathbb{R}}_{+}^{n}$. Assume that $\deg g=d$. Denote by $g^{\infty}$ the leading term of $g$, meaning that $g_{i}^{\infty}$ is the homogenous term of degree $d$ of $g_{i}$, for $i\in[n]$.
Recall from (gowda2016polynomial, , Corollary 7.2) that if $f$ or $f^{\infty}$ is copositive and $\operatorname{Sol}(f^{\infty})=\{0\}$, then $\operatorname{Sol}(f+a)$ is nonempty and compact for every vector $a$ in $\operatorname{\mathbb{R}}^{n}$.
Proposition 4
Suppose that $f$ or $f^{\infty}$ is copositive and $\operatorname{Sol}(f^{\infty})=\{0\}$. The following statements hold:
  • $\rm a)$
    There exists a generic set $\mathcal{O}$ in $\operatorname{\mathbb{R}}^{n}$ such that $\operatorname{CP}(f+a)$ is $\operatorname{\mathcal{D}}_{0}$ for every $a\in\mathcal{O}$.
  • $\rm b)$
    For every $\mu>0$, there exists $\varepsilon>0$ such that $\operatorname{Sol}(f+a)$ is nonempty and $\operatorname{Sol}(f+a)\subseteq\operatorname{Sol}(f)+\mathbb{B}(0,\mu)$ for any $a\in\operatorname{\mathbb{R}}^{n}$ satisfying $\|a\|<\varepsilon$.
Proof
The proof of the first statement is analogous to the proof of Proposition 2 shown in the appendix. The second one is obtained by applying directly (gowda2016polynomial, , Corollary 7.2) and (hieu2020solution, , Corollary 4.1) to polynomial complementarity problems. ∎
It is noteworthy that Proposition 4 enables us to solve approximately a PCP satisfying the assumptions therein. We take $a\in\operatorname{\mathbb{R}}^{n}$ randomly and $\|a\|$ small enough, and then solve $\operatorname{CP}(f+a)$ by using the method proposed in Section 6.1. A solution of $\operatorname{CP}(f+a)$ can be considered as an approximate solution to the original problem.
Example 3
Consider the polynomials in two variables $x_{1},x_{2}$ given in (hieu2019r0, , Example 5.3), $f_{1}=f_{2}=x_{1}^{2}+x_{2}^{2}-1$. It is easy to check that $f_{1}^{\infty}=f_{2}^{\infty}=x_{1}^{2}+x_{2}^{2}$ and $f^{\infty}$ is copositive and that $\operatorname{Sol}(f^{\infty})=\{0\}$. Moreover, $\operatorname{CP}(f)$ has infinitely many solutions
$\operatorname{Sol}(f)=\{(\alpha_{1},\alpha_{2})\in\operatorname{\mathbb{R}}^{2}:\alpha_{1}^{2}+\alpha_{2}^{2}=1,\alpha_{1}\geq 0,\alpha_{2}\geq 0\}.$
We observer that $\operatorname{CP}(f+a)$, where $a=(a_{1},a_{2})$, is $\operatorname{\mathcal{D}}_{0}$ if any only if $a_{1}\neq a_{2}$. Taking $a_{1}=1/10^{6}$ and $a_{1}=1/10^{5}$ and computing solutions to $\operatorname{CP}(f+a)$ by using the proposed method, the perturbed problem has unique solution $\alpha_{a}=(0.999999,0.000000)$ which is very close to a solution $(1,0)$ of $\operatorname{CP}(f)$.
Appendix
The notions of differentiable maps and regular values can be found in monographs on differential geometry, for example, loring2011introduction .
Proof of Proposition 2. Let $\ell$ be a subset of $[n]$ and $\mathcal{V}_{\ell}$ be the set of all $\alpha\in\operatorname{\mathbb{C}}^{n}$ such that $\alpha_{i}\neq 0$ if $i\in\ell$ and $\alpha_{i}=0$ if $i\notin\ell$. Clearly, $\mathcal{V}_{\emptyset}=\{0\}$ and $\operatorname{\mathbb{C}}^{n}$ is the disjoint union of $\mathcal{V}_{\ell}$ for $\ell\subset[n]$.
Suppose that $\ell\neq\emptyset$ and consider the differentiable map $\Phi_{\ell}:\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{(\rho-1)\times n}\times\mathcal{V}_{\ell}\to\operatorname{\mathbb{C}}^{\#\ell}$ defined by
$\Phi_{\ell}(a,Q,x_{\ell})=\left(a_{i}x_{i}+Q_{i}X^{T}\right)_{i\in\ell},$
where $Q\in\operatorname{\mathbb{R}}^{(\rho-1)\times n}$ and $X$ is given in (4.1). Obviously, its Jacobian matrix with respect to $(a,Q,x_{\ell})$, denoted by $J_{\ell}$, has size $(n\cdot\rho\cdot\#\ell)\times\#\ell$. Moreover, $J_{\ell}$ has a diagonal submatrix of size $\#\ell\times\#\ell$, and the entries on the diagonal are the coordinates of $x_{\ell}$. Since all coordinates of $x_{\ell}$ are non-zero, we conclude that $J_{\ell}$ is full rank. It follows that the origin 00 in $\operatorname{\mathbb{C}}^{\#\ell}$ is a regular value of $\Phi_{\ell}$. By the Sard Theorem with parameter (dang2016well, , Theorem 2.4), there exists a generic set $\mathcal{O}_{\ell}$ in $\operatorname{\mathbb{R}}^{n\times\rho}$ such that for every $(a,Q)$ in $\mathcal{O}_{\ell}$, 00 in $\operatorname{\mathbb{C}}^{\#\ell}$ is a regular value of $\Phi_{\ell,a,Q}:\mathcal{V}_{\ell}\to\operatorname{\mathbb{C}}^{\#\ell}$ defined by $\Phi_{\ell,a,Q}(x)=\Phi_{\ell}(a,Q,x)$. It is clear that
$\Phi_{\ell,a,Q}^{-1}(0)=\{x_{\ell}:a_{i}x_{i}+Q_{i}X^{T}=0,i\in\ell\},$
and if this set is non-empty then it has finitely many points (loring2011introduction, , Theorem 9.9).
Furthermore, we observe that
$V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\subseteq\{0\}\cup\big{(}\cup_{\ell\subseteq[n],\ell\neq\emptyset}\Phi_{\ell,a,Q}^{-1}(0)\big{)},$
and that the finite union $\mathcal{O}:=\bigcup_{\ell\subseteq[n],\ell\neq\emptyset}\mathcal{O}_{\ell}$ is generic in $\operatorname{\mathbb{R}}^{n\times\rho}$. Hence, for every $(a,Q)$ in $\mathcal{O}$, $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ has finitely many points. ∎